OpenStudy (anonymous):

I want a elegant method to solve this question i have already solved it though.

OpenStudy (anonymous):

The sum of three numbers in an G.P is 70 , the two extremes be multiplied each by 4 and the mean by 5, the products are in A.P . Find the numbers

OpenStudy (anonymous):

I am getting two values of "r" the common differnce

Miracrown (miracrown):

I need to see the original wording

OpenStudy (anonymous):

wording? that is the exact question as from the textbook

Miracrown (miracrown):

the question is not clearly stated

OpenStudy (anonymous):

let the final numbers in AP be a-d,a,a+d then (a-d)/4 ,a/5 ,(a+d)/4 add upto 70

OpenStudy (anonymous):

and these are in GP too

OpenStudy (anonymous):

Well i considered a/r , a , ar

OpenStudy (anonymous):

so a/2 +a/5 =70 so a= 100

OpenStudy (anonymous):

but the mean of the GP is 20

OpenStudy (anonymous):

a +ar + ar^2 =70r a(1+r+r^2) =70r From the second piece of information I got 4r^2 -10r +4 =0 and thus r = 1/2 and 2

OpenStudy (anonymous):

I don't actually need extra large explanation and information , i will understand and percieve no problem

Miracrown (miracrown):

  10, 20 and 40